OpenStudy (anonymous):

Create your own radical expression, containing a numerical coefficient and two different variables, which can be simplified. Explain, in complete sentences, how you developed the expression, the process of simplifying it, and provide the final simplified expression.

OpenStudy (a_clan):

\[5\sqrt{x}+y \sqrt{x}\] contains a numerical coefficient, two variables and square root. In order to simplify it, we take square_root(x) common. Or in other words ,we use distributive property to get the following simplified expression. \[(5+y)\sqrt{x}\]